# Provably Unlearnable Examples

### v1

@@ 7.2. Robustness of PUE against GAs
In this section, we would like to understand how robust PUEs are when the classifier trained on them has a very limited probability of being sampled into the certified parameter set $\hat{\Theta}$. This setting aligns with real-world cases in which unauthorized classifiers are distinct from the surrogates used in the certification.
@@
We particularly focus on the robustness of UEs by showing some hardness results of recovery attacks against classifiers trained on PUE, EMN, and OPS, following common training routines. Furthermore, we evaluate the performance of PUE in defending against various training approaches to verify its practicality. Specifically, we compare the empirical clean test accuracy of unauthorized classifiers trained on PUE, EMN and OPS using four representative training strategies (i.e., MixUp, CutOut, Fast Autoaugment, and Adversarial Training). The results are deferred to Appendix C.3.
Figure 6. The robustness of PUE and OPS against recovery attacks.
@@
The comparison between EMN and PUE is presented in Figure 5. Compared to the results in Figure 4, the clean test accuracy of the normally trained classifiers can be better restored. On CIFAR10, the poisoned classifiers of PUE and EMN have initial accuracy scores of $11.62\%$ and $11.89\%$, respectively. After the recovery, the EMN classifier has $78.55\%$ accuracy and the PUE one has $72.37\%$ accuracy. On the side of the best possible accuracy, EMN can achieve $100\%$ and PUE reaches $99.55\%$. Similarly, the EMN classifier can be restored to $43.01\%$ accuracy on CIFAR100, while the PUE classifier accuracy achieves $32.19\%$. However, the classifier trained on PUE demonstrates better robustness than the EMN classifier in both the best possible case and the recovery attack case. Most accuracy can be recovered within an $\eta$ of $1.0$, especially for CIFAR10. This is also why we certify at most an $\eta$ of $1.0$ in Section 7.1.
@@
Aside from the above comparison, we notice several intriguing phenomena of PUE and OPS in Figure 6. OPS is a surrogate-free PAP noise. It searches for the pixel leading to the most significant change in the input data distribution, for each class in the dataset. Therefore, OPS is supposed to have higher robustness since it is model-agnostic and can be generalized to different classifiers based on experiments. It can be observed from the figure that, though OPS has a lower empirically recovered accuracy than PUE, its best recoverable accuracy gets higher than that of PUE rapidly when $\eta$ grows. This implies that OPS has remarkable local robustness, possibly due to the effective manipulation of the input data distribution, which makes it hard to find proper clean training samples to reduce the clean generalization error of the trained classifier. However, when a set of plausible training samples is selected (e.g., the Best from OPS case), OPS can be more vulnerable than other PAPs. Specifically, PUEs reduce at most $54.4\%$ of the accuracy on CIFAR100. The weakness becomes more obvious when the perturbations/uncertainty in the weights increase beyond a certain threshold.
@@
The comparisons suggest that training GAs on PUE can lead to more robust poisoning results, especially against best-case recovery attacks. We attribute this to two reasons. First, theoretically, $\hat{\Theta}$ occupies a non-zero probability measure on every infinitesimal region on $\mathbb{R}^{d}$, the certified accuracy upper bound holds for all possible classifiers with parameters in $\mathbb{R}^{d}$, with probability at least $q$. Therefore, these empirically trained classifiers have non-zero probabilities of being included in $\hat{\Theta}$. In practice, a larger certified parametric radius $\eta$ helps increase the probability measure at classifiers far from the surrogate. Second, From the perspective of uncertainty, the loss function 13 encourages PUE to reduce the accuracy of a collection of classifiers with Gaussian-modeled uncertainty in their parameters. Consequently, the collective robustness of PUE towards classifiers showing uncertainties in their parameters is enhanced.
Figure 7. The $(q,\eta)$-Learnability scores under different values of $q$ are plotted on three data distributions (i.e., CIFAR10, CIFAR100, and ImageNet). At the same $\eta$, a greater probability $q$ can certify a higher $(q,\eta)$-Learnability.
7.3. Ablation Study
@@
We conduct an ablation study on the trade-off between the standard deviation $\sigma$ of the smoothing noise and the learnability scores over different factors. First, we evaluate how the train-time noise during PUE surrogate optimization affects the trade-off. We train PUE-B surrogates on the CIFAR10/CIFAR100 datasets under different train-time noises and measure their certified learnability under smoothing noises with various $\sigma$ values. We fix the value of $q$ at $0.9$ in all cases. The results on CIFAR10 and CIFAR100 are in Table 7 and Table 8 of Appendix C.2, respectively. The certification returns Abstain when a valid binomial confidence interval cannot be found. The abstentions are marked by “--” in the tables. The highest certified learnability under each $\eta$ is highlighted.
@@
The numbers in Table 7 provide some intriguing observations about the trade-off. First, the largest certifiable parametric radius $\eta$ positively correlates with $\sigma$. However, large $\sigma$ values result in decreases in the $(q,\eta)$-Learnability scores at small certified radii, which indicates that the negative impact of the smoothing noise towards the tightness gap is significant in these cases. On the train-time noise side, generally, greater train-time noises make the resulting surrogates more endurable to large smoothing noises in the certification. Thus, the certification performs better in large-$\sigma$ cases and produces better average certified learnability over different $\eta$ values. Generally, when the train-time noise matches the $\sigma$ value, the highest certified learnability scores can be found. For example, certifying using $\sigma=0.25$ obtained the highest $(q,\eta)$-Learnability score when the standard deviation of the train-time noise is $0.25$. We observe similar trends in the CIFAR100 certification. This phenomenon suggests that the train-time noise should match the parametric smoothing noise to reduce the tightness gap in the certification.
@@
Next, we evaluated the $(q,\eta)$-Learnability under different $q$ values to see how the certified learnability can be affected. We select $0.25$ as the standard deviation for the train time noise and the parametric smoothing noise in this study. The learnability scores are illustrated in Figure 7. From the figure, certifying at a higher probability $q$ compromises the maximal certifiable parametric radius across all datasets. However, the learnability scores certified at the same radius with larger $q$ values are usually larger than those certified with small $q$ values, which reveals tighter guarantees on the best cases unauthorized classifiers can achieve. At smaller parametric radii, the certified learnability is less distinguishable among different $q$ values. Nevertheless, the discrepancy becomes obvious when $\eta$ grows. In this paper, to ensure the $(q,\eta)$-Learnability can be certified at reasonable $\eta$ values and be held with a sound probability, we select $q=0.9$ in the certification experiments.
We also tested the effective noise level for training the surrogate and PUEs. We found that the surrogate and PUEs could converge rapidly on a Gaussian noise whose $\sigma^{\prime}\leq 0.25$. However, when the noise further increases to $\sigma^{\prime}\geq 0.3$, the training fails to converge in a limited time. More experiments on the surrogate architectures and training cost are in Appendix C.
8. Discussion and Limitation
In the following paragraphs, we will discuss the limitations of our methods and some possible improvements that can be made in the future.
@@ 8. Discussion and Limitation
Certified parameter set.  A crucial step for improving the certificate is to increase the certified parametric radius $\eta$ such that the certificate can cover unauthorized classifiers with more diverse parameters. Based on the theory and the experiments, we can find a negative correlation between $q$ and the maximum of certifiable $\eta$. This trade-off suggests that confident certifications can only be made at small $\eta$. While thoroughly breaking this trade-off seems impossible, it is possible to improve $\eta$ by meticulously designing the parametric smoothing noise and the surrogate training method. In Theorem 5.2, the certification abstains if $\overline{q}$ is too large such that a valid Binomial confidence interval cannot be found. To avoid the abstention at large $\eta$, we should use a substantial $\sigma$ since $\eta$ scales with $\sigma$. Next, to make such a substantial $\sigma$ functioning, we should have a surrogate classifier that can tolerate a significant level of smoothing noise while not degrading the tightness of the certification. Through experiments, we verify that larger parametric smoothing noise indeed leads to a broader certifiable parametric radius. Nevertheless, increasing the parametric radius $\eta$ cannot be solely resorted to enlarging the smoothing noise since larger noise can also compromise the tightness of the $(q,\eta)$-Learnability. Specifically, too-large noises negatively impact the testing accuracy of the classifiers randomized from an insufficiently trained surrogate and make them fail to obtain plausible clean test accuracy scores. We propose a baseline solution in this paper to alleviate the negative effects of the parametric noise, but we reckon further endeavors can be made to increase $\eta$ at high values of $q$.
@@
Tightness gap of the certification.  Given the noticeable connection between certified learnability and certified robustness, a possible way to further improve the tightness of the certification might be using a combination of adversarial training techniques, such as AWP (Wu et al., 2020) and SmoothAdv(Salman et al., 2019), to find the most critical perturbation direction for the weights during training the surrogate. In contrast, optimizing the PAP noise against these carefully augmented surrogates may lead to more robust UEs. Another possible pathway is to select better parametric smoothing noise. Though this simple uni-variance Gaussian noise used in this paper endorses a tidy form of $(q,\eta)$-Learnability, we suspect the tightness gap can be further reduced by introducing better smoothing noises. First, the probability density of Gaussian might not be optimal in discovering the best cases of classifiers. Intuitively, the optimal parametric smoothing noise should have a large probability of drawing parameters that can lead to the highest clean test accuracy. Second, the weights in different layers of a classifier may respond differently against the noise added to them, rendering uni-variance noises less efficient in surrogate augmentation and certification. A better option could be applying different levels of train-time noise in different layers to cope with the layer-wise noise sensitivities and using corresponding parametric smoothing noise in the certification.
9. Related work
@@
Perturbative availability poison and UEs. PAPs emerged as a response to unauthorized models trained to infer private datasets. There were early attempts that generate perturbations dynamically on surrogate models, which is computationally intensive (Shen et al., 2019; Feng et al., 2019). To alleviate the cost, error-minimizing perturbations were introduced to produce unlearnable data samples (Huang et al., 2021). Subsequently, the training cost of the perturbation was further reduced by using pretrained surrogate classifiers or generalized neural tangent kernels (Tao et al., 2021; Fowl et al., 2021; Yuan and Wu, 2021; Chen et al., 2023). A series of works also resort to surrogate-free perturbations (Evtimov et al., 2021; Yu et al., 2022; Sandoval-Segura et al., 2022; Wu et al., 2023). In addition, enhancements were proposed to make transferable PAP noises or robust UEs against adversarial training  (Fu et al., 2022; Wen et al., 2022; Ren et al., 2022). We also notice some work making efforts to address the uncertainty of UEs (Lu et al., 2023; He et al., 2023). Nevertheless, current methods fall short of offering a provable guarantee on the robustness of UEs against uncertainties in learning algorithms and adaptive adversaries in the field.
@@
Certified robustness via randomized smoothing.  Following previous certified robustness works using differential privacy and Rényi divergence (Lecuyer et al., 2019; Li et al., 2019), randomized smoothing was first proposed by Cohen et al. as a tool for supplying a tight $\ell_{2}$ robustness certificate for black-box functions, based on Neyman-Pearson Lemma (NPL) (Cohen et al., 2019). Though randomized smoothing places no restriction on its based classifiers, a line of work aimed at improving the base classifiers to obtain a better trade-off between the certified accuracy and the certified radius (Salman et al., 2019; Zhai et al., 2020; Jeong and Shin, 2020; Jeong et al., 2021, 2023). On the other hand, some trials have been made to design more capable smoothing noises, smoothing pipelines, and certification algorithms beyond NPL (Fischer et al., 2020; Li et al., 2021; Hao et al., 2022; sukenik2022intriguing; Cullen et al., 2022; Dvijotham et al., 2020; Zhang et al., 2020; Salman et al., 2020). Intriguingly, a series of works discovered that the maximum of certifiable $\ell_{p}$ ($p>2$) radius shrinks with the increasing input dimensionality (Kumar et al., 2020; Yang et al., 2020; Li et al., 2022). Randomized smoothing has also been applied to the training phase(Zhang et al., 2022; Weber et al., 2022; Kumar et al., 2023; Zhang et al., 2023). Last but not least, randomized smoothing has been extended to protect tasks such as object detection (Chiang et al., 2020) and watermark verification (Bansal et al., 2022). The works in this area focus on the certified robustness based on worst-case models rather than examining the certified learnability of datasets through the analysis of the best-case models trained on them.
10. Conclusion
@@
In this paper, we propose a certification mechanism for deriving the possibly best clean data utility that unauthorized classifiers trained on UEs can achieve. The certified $(q,\eta)$-Learnability indicates that, with a probability at least $q$, the unauthorized classifiers whose weights are from a specific parameter subspace have a guaranteed upper bound on their clean test accuracy. Moreover, we propose a way to craft provably unlearnable examples that obtain smaller $(q,\eta)$-Learnability scores, compared with the ones crafted by existing methods. Our certification mechanism takes the first step towards certifiably robust and provably effective UEs, which can provide a rigid guarantee on the protection level of PAP noises for data availability control. However, there exists a tightness gap between the certified $(q,\eta)$-Learnability and True Learnability. Such a tightness gap can be better alleviated in the future to make the certification more tenable. Crucially, such alleviation may also expand the space of certifiable parameters. Finally, we reckon that future attempts can be made to extend PUE to sample-wise noises and reduce the training cost of making PUEs.
11. Acknowledgment

### v2

@@ VII-B Robustness of PUE against GAs
In this section, we aim to understand how robust PUEs are when the classifier trained on them is outside of the certified parameter set $\hat{\Theta}$ or evades the certified learnability scores with a probability of $1-q$. This setting aligns with real-world cases in which unauthorized classifiers are distinct from the surrogates used in the certification.
@@
We particularly focus on the robustness of UEs by showing some hardness results of recovery attacks against classifiers trained on PUE, EMN, and OPS, following common training routines. Furthermore, we evaluate the performance of PUE in defending against various training approaches to verify its practicality. Specifically, we compare the empirical clean test accuracy of unauthorized classifiers trained on PUE, EMN and OPS using four representative training strategies (i.e., MixUp, CutOut, Fast Autoaugment, and Adversarial Training). The results are deferred to Appendix -D.
Figure 6: The robustness of PUE and OPS against recovery attacks.
Figure 7: The $(q,\eta)$-Learnability scores under different values of $q$ are plotted on three data distributions (i.e., CIFAR10, CIFAR100, and ImageNet). At the same $\eta$, a greater probability $q$ can certify a higher $(q,\eta)$-Learnability.
@@
The comparison between EMN and PUE is presented in Figure 5. Compared to the results in Figure 4, the clean test accuracy of the normally trained classifiers can be better restored. On CIFAR10, the poisoned classifiers of PUE and EMN have initial accuracy scores of $11.62\%$ and $11.89\%$, respectively. After the recovery, the EMN classifier has $78.55\%$ accuracy and the PUE one has $72.37\%$ accuracy. On the side of the best possible accuracy, EMN can achieve $100\%$ and PUE reaches $99.55\%$. Similarly, the EMN classifier can be restored to $43.01\%$ accuracy on CIFAR100, while the PUE classifier accuracy achieves $32.19\%$. However, the classifier trained on PUE demonstrates better robustness than the EMN classifier in both the best possible case and the recovery attack case. Most accuracy can be recovered within an $\eta$ of $1.0$, especially for CIFAR10. This is also why we certify at most an $\eta$ of $1.0$ in Section VII-A.
@@
Aside from the above comparison, we notice several intriguing phenomena of PUE and OPS in Figure 6. OPS is a surrogate-free PAP noise. It searches for the pixel leading to the most significant change in the input data distribution, for each class in the dataset. Therefore, OPS is supposed to have higher robustness since it is model-agnostic and can be generalized to different classifiers based on experiments. It can be observed from the figure that, though OPS has a lower empirically recovered accuracy than PUE, its best recoverable accuracy gets higher than that of PUE rapidly when $\eta$ grows. This implies that OPS has remarkable local robustness, possibly due to the effective manipulation of the input data distribution, which makes it hard to find proper clean training samples to reduce the clean generalization error of the trained classifier. However, when a set of plausible training samples is selected (e.g., the Best from OPS case) for the recovery attack, OPS can be more vulnerable than other PAPs. Specifically, PUEs reduce at most $54.4\%$ of the accuracy on CIFAR100. The weakness becomes more obvious when the perturbations/uncertainty in the weights increase beyond a certain threshold.
@@
The comparisons suggest that training GAs on PUE can lead to more robust poisoning results, especially against best-case recovery attacks. We attribute this to two reasons. First, theoretically, $\hat{\Theta}$ occupies a non-zero probability measure on every infinitesimal region on $\mathbb{R}^{d}$. Therefore, these empirically trained classifiers have non-zero probabilities of being included in $\hat{\Theta}$. In practice, a larger certified parametric radius $\eta$ helps increase the probability measure at classifiers far from the surrogate. Second, the loss function 13 encourages PUE to reduce the accuracy of a collection of classifiers with Gaussian-modeled uncertainty in their parameters. Consequently, the collective robustness of PUE towards classifiers showing uncertainties in their parameters is enhanced.
VII-C Ablation Study
@@
We conduct an ablation study on the trade-off between the STD $\sigma$ of the smoothing noise and the learnability scores over different factors. First, we evaluate how the train-time noise during PUE surrogate optimization affects the trade-off. We train PUE-B surrogates on the CIFAR10/CIFAR100 datasets under different train-time noises and measure their certified learnability under smoothing noises with various $\sigma$ values. We fix the value of $q$ at $0.9$ in all cases. The results on CIFAR10 and CIFAR100 are in Table VII and Table VIII of Appendix -C, respectively. The certification returns Abstain when a valid binomial confidence interval cannot be found. The abstentions are marked by “--” in the tables. The highest certified learnability under each $\eta$ is highlighted.
@@
The numbers in Table VII provide some intriguing observations about the trade-off. First, the largest certifiable parametric radius $\eta$ positively correlates with $\sigma$. However, large $\sigma$ values result in decreases in the $(q,\eta)$-Learnability scores at small certified radii, which indicates that the negative impact of the smoothing noise towards the tightness gap is significant in these cases. On the train-time noise side, generally, greater train-time noises make the resulting surrogates more endurable to large smoothing noises in the certification. Thus, the certification performs better in large-$\sigma$ cases and produces better average certified learnability over different $\eta$ values. Generally, when the train-time noise matches the $\sigma$ value, the highest certified learnability scores can be found. For example, certifying using $\sigma=0.25$ obtained the highest $(q,\eta)$-Learnability score when the STD of the train-time noise is $0.25$. We observe similar trends in the CIFAR100 certification. This phenomenon suggests that the train-time noise should match the parametric smoothing noise to reduce the tightness gap.
@@
Next, we evaluated the $(q,\eta)$-Learnability under different $q$ values to see how the certified learnability can be affected. We select $0.25$ as the STD for the train time noise and the parametric smoothing noise in this study. The learnability scores are illustrated in Figure 7. From the figure, certifying at a higher probability $q$ compromises the maximal certifiable parametric radius across all datasets. However, the learnability scores certified at the same radius with larger $q$ values are usually larger than those certified with small $q$ values, which reveals tighter guarantees on the best cases unauthorized classifiers can achieve. At smaller parametric radii, the certified learnability is less distinguishable among different $q$ values. Nevertheless, the discrepancy becomes obvious when $\eta$ grows. In this paper, to ensure the $(q,\eta)$-Learnability can be certified at reasonable $\eta$ values and be held with a sound probability, we select $q=0.9$ in the certification experiments.
VIII Discussion and Limitation
In the following part, we will discuss the limitations of our methods and explore potential avenues for improvement.
@@ VIII Discussion and Limitation
Certified parameter set.  A crucial step for improving the certificate is to increase the certified parametric radius $\eta$ such that the certificate can cover unauthorized classifiers with more diverse parameters. Based on the theory and the experiments, we can find a negative correlation between $q$ and the maximum of certifiable $\eta$. This trade-off suggests that confident certifications can only be made at small $\eta$. While thoroughly breaking this trade-off seems impossible, it is possible to improve $\eta$ by meticulously designing the parametric smoothing noise and the surrogate training method. In Theorem 1, the certification abstains if $\overline{q}$ is too large such that a valid Binomial confidence interval cannot be found. To avoid the abstention at large $\eta$, we should use a substantial $\sigma$ since $\eta$ scales with $\sigma$. Next, to make such a substantial $\sigma$ functioning, we should have a surrogate classifier that can tolerate a significant level of smoothing noise while not degrading the tightness of the certification. Through experiments, we verify that larger parametric smoothing noise indeed leads to a broader certifiable parametric radius. Nevertheless, increasing the parametric radius $\eta$ cannot be solely resorted to enlarging the smoothing noise since larger noise can also compromise the tightness of the $(q,\eta)$-Learnability. Specifically, too-large noises negatively impact the testing accuracy of the classifiers randomized from an insufficiently trained surrogate and make them fail to obtain plausible clean test accuracy scores. We propose a baseline solution in this paper to alleviate the negative effects of the parametric noise, but we reckon further endeavors can be made to increase $\eta$ at high values of $q$.
@@
Tightness gap of the certification.  Given the noticeable connection between certified learnability and certified robustness, a possible way to further improve the tightness of the certification might be using a combination of adversarial training techniques, such as AWP [36] and SmoothAdv[37], to find the most critical perturbation direction for the weights while training the surrogate. In contrast, optimizing the PAP noise against these carefully augmented surrogates may lead to more robust UEs. Another possible pathway is to select better parametric smoothing noise. Though this simple uni-variance Gaussian noise used in this paper endorses a tidy form of $(q,\eta)$-Learnability, we suspect the tightness gap can be further reduced by introducing better smoothing noises. First, the probability density of Gaussian might not be optimal in discovering the best cases of classifiers. Intuitively, the optimal parametric smoothing noise should have a large probability of drawing parameters that can lead to the highest clean test accuracy. Second, the weights in different layers of a classifier may respond differently against the noise added to them, rendering uni-variance noises less efficient in surrogate augmentation and certification. A better option could be applying various levels of train-time noise in different layers to cope with the layer-wise noise sensitivities and using corresponding parametric smoothing noise in the certification.
Real-world practicality.  Certified $(q,\eta)$-Learnability is a versatile framework that can be applied to different data modalities and tasks, as the certification holds in the parameter space and $A_{{\mathbb{D}}}(\hat{\theta})$ can represent any function that maps model parameters to a scalar value. As a proof of concept, we certify the unlearnable SST2 training set with Error-min-0 modifications [38] via an LSTM surrogate. A $(0.9,\eta)$-learnability of $0.61$ can be certified at $\eta=1$ while the accuracy at $\eta=0$ is $0.52$. Furthermore, the certification framework and PUE generation process may introduce additional computational overhead. We benchmark this overhead across various model architectures and datasets, as shown in Tables XIII and XIV. The overhead is confined to the certification and PUE generation stages and remains manageable. To reduce the cost, techniques such as perturbing only specific layers during PUE generation can be explored. Importantly, the application of PUE does not impact real-time data streaming, as PAP noise for each data category can be pre-generated and applied to corresponding streaming data. Finally, PUE requires a high perturbation rate, which means that a small proportion of PUEs in a clean dataset will not hinder the training process for authorized users. Legitimate users can also be provided with the PUE perturbations to restore data learnability. These features suggest that PUE is compatible with other defenses, such as model watermarking and unlearning, as the functionality of the watermark or retained set can be maintained.
IX Related work
@@
Perturbative availability poison and UEs. PAPs emerged as a response to unauthorized models trained to infer private datasets. There were early attempts that generate perturbations dynamically on surrogate models, which is computationally intensive [39, 10]. To alleviate the cost, error-minimizing perturbations were introduced to produce unlearnable data samples [9]. Subsequently, the training cost of the perturbation was further reduced by using pretrained surrogate classifiers or generalized neural tangent kernels [11, 8, 14, 15]. A series of works also resort to surrogate-free perturbations [40, 41, 42, 35]. In addition, enhancements were proposed to make transferable PAP noises or robust UEs against adversarial training  [13, 43, 44]. We also notice some work making efforts to address the uncertainty of UEs [45, 16]. Particularly, the sharpness-aware method SAPA [16] leverages weight perturbations to approximate the worst-case model for the poison attack. However, the perturbation considered in SAPA is model- and data-dependent and may not be effective with variations in the type of loss function. Moreover, all the current methods fall short of offering a provable guarantee on the robustness of UEs against uncertainties in learning algorithms and adaptive adversaries in the field.
@@
Certified robustness via randomized smoothing.  Following previous certified robustness works using differential privacy and Rényi divergence [46, 47], randomized smoothing was first proposed by Cohen et al. as a tool for supplying a tight $\ell_{2}$ robustness certificate for black-box functions, based on Neyman-Pearson Lemma (NPL) [21]. Though randomized smoothing places no restriction on its based classifiers, a line of work aimed at improving the base classifiers to obtain a better trade-off between the certified accuracy and the certified radius [37, 48, 49, 50, 51]. On the other hand, some trials have been made to design more capable smoothing noises, smoothing pipelines, and certification algorithms beyond NPL [22, 23, 24, 25, 26, 27, 28, 29]. Intriguingly, a series of works discovered that the maximum of certifiable $\ell_{p}$ ($p>2$) radius shrinks with the increasing input dimensionality [52, 30, 53, 54]. Moreover, randomized smoothing has been extended to poisoning defenses[55, 56, 57, 58], object detection [32], and watermark verification [33]. The works in this area focus on the certified robustness based on worst-case models rather than examining the certified learnability of datasets through the analysis of the best-case models trained on them.
X Conclusion
@@
In this paper, we propose a certification mechanism for deriving the possibly best clean data utility that unauthorized classifiers trained on UEs can achieve. The certified $(q,\eta)$-Learnability indicates that, with probability at least $q$, the unauthorized classifiers whose weights are from a specific parameter subspace have a guaranteed upper bound on their clean test accuracy. Moreover, we propose a way to craft provably unlearnable examples that obtain smaller $(q,\eta)$-Learnability scores, compared with the ones crafted by existing methods. Our certification mechanism takes the first step towards certifiably robust and provably effective UEs, which can provide a rigid guarantee on the protection level of PAP noises for data availability control. However, there exists a tightness gap between the certified $(q,\eta)$-Learnability and True Learnability. Such a tightness gap can be better alleviated in the future to make the certification more tenable. Crucially, such alleviation may also expand the space of certifiable parameters to cover greater parameter variance resulting from training techniques such as adversarial training. Finally, we reckon that future attempts can be made to extend PUE to sample-wise noises and reduce the training cost of making PUEs.
Acknowledgments